On the x axis there are point charges 3μC and 5μC located at the positions x=-2cm and x=3cm respectively. Find the direction and magnitude of the electric field at the origin.


1
Expert's answer
2021-04-23T11:10:00-0400
"E=9\cdot10^9\frac{3\cdot10^{-6}}{0.02}-9\cdot10^9\frac{5\cdot10^{-6}}{0.03}=-150000\frac{V}{m}"

Thus, it has negative x-direction.